Hello!
A machine can fill 24 boxes in one hour. At this rate, how many boxes can it fill in five minutes?
Thank you!

Answer by jim_thompson5910(35256)

1 hour = 60 minutes


24 boxes/60 minutes = x boxes/5 minutes

24/60 = x/5

24*5 = 60*x

120 = 60x

60x = 120

x = 120/60

x = 2


So the machine can fill 2 boxes in 5 minutes.


Let me know if you need more help or if you need me to explain a step in more detail.
Feel free to email me at jim_thompson5910@hotmail.com
or you can visit my website here: http://www.freewebs.com/jimthompson5910/home.html

Thanks,

Jim